OpenStudy (anonymous):

Please help as soon as possible. I really need this explained to me. Just a few questions, also some are answered i just do not know if it is correct. ☺ Thanks so much! 2. Given the function k(x) = x2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(x + c) k(x) + c k(cx) c • k(x)

OpenStudy (anonymous):

One that i answered.. 1. Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. f(x) = (x -2)(x + 2) The function g(x) is a sum of squares. g(x) = (x - 7)^2 The function h(x) is a perfect square trinomial. h(x) = x^2 + 25x + 144 The function j(x) can only have a GCF factored out of it. j(x) = 6x^2 + 4x + 8 lol i am really..really bad at math..please don't judge! ^.^

OpenStudy (jack1):

just remember FOIL methos so multiply: First Terms (x -2)(x + 2) first terms are x and x = x^2 Outer terms so outer terms are x and 2 = 2x Inner Terms inner terms are -2 and x = -2x Last Terms last terms are -2 and 2 = -4 now add all these multiples together to get your quadratic = (x -2)(x + 2) = x^2 + 2x - 2x - 2 = x^2 - 2

OpenStudy (jack1):

that's how to get the quadratic, as you already have the factors [ which are (x -2)(x + 2) ]

OpenStudy (anonymous):

Thank you! This helped a lot. ☺ ♥ ☺ Could you help with the first one? How would i do that??

OpenStudy (jack1):

sorry, which one...?

OpenStudy (anonymous):

It's ok :) i was talking about the first one. Then i looked at your response again! I got it ^.^ Thanks lots ♥

OpenStudy (jack1):

all good, np ;)

OpenStudy (anonymous):

Yupp ^.^